QUANTUM MONTE CARLO STUDIES OF DOPED SOLID PARAHYDROGEN: STRUCTURE AND IR ABSORPTION SPECTRA

R. J. HINDE, Department of Chemistry, University of Tennessee, Knoxville, TN 37996.

The infrared absorption spectrum of atom- and molecule-doped solid parahydrogen (p-H_{2}) exhibits features near 4160 cm^{−1} which are absent from the spectrum of pure solid p-H_{2}. These absorption features are associated with the Q_{1}(0) pure vibrational transitions of H_{2} molecules located near impurities; the transition moment responsible for this infrared activity arises from weak overlap-induced dopant–H_{2} and H_{2}–H_{2} dipoles, and from dopant-induced symmetry breaking of the p-H_{2} crystal lattice. Because the transition dipole moment for these infrared features depends sensitively on the structure of the p-H_{2} matrix surrounding the dopant, studies of the Q_{1}(0) absorption feature could provide detailed information about the physics of solvation in highly quantum systems. We present the results of diffusion quantum Monte Carlo (DQMC) simulations of the Q_{1}(0) absorption lineshape for isolated dopants in solid p-H_{2}, and briefly discuss the structural and dynamical information provided by our DQMC simulations.